To highlight relevant issues that are still relevant I have moved old/historical information to the end.
Update September 24th, 2020
- GPI is not available in semester 20B or later due to being removed from science operations for the GPI 2.0 upgrade and move to Gemini North.
Update March 19th, 2020
- GPI is now warming up since Monday the 16th due to the unplanned shutdown due to the Covid-19 outbreak. The instrument is now safe in a shutdown state on the telescope.
Update March 3rd, 2020
- GPI is not offered in Semester 2020B as the instrument will leave GS in early 20B for upgrades and later commissioning at GN.
Update June 26th, 2019
As a friendly reminder proposers should be aware that:
All K2 observations (see on on page http://www.gemini.edu/instrumentation/gpi/capability#Configuration) are taken without any evaluation contrast and data-quality is only based on DIMM/MASS measurements indicating that the data is taken in the required IQ constraint.
All NRM data are only guaranteed to be taken under the requested conditions as any other performance indicators are not available, due to no pipeline reduction of NRM data. See the comment above the second table on page http://www.gemini.edu/instrumentation/gpi/capability#Configuration.
Data taken in any mode with Zenith Distance larger than 40 degrees are only evaluated based on DIMM/MASS measurements indicating that the data is taken in the required IQ constraint. On page https://www.gemini.edu/instrumentation/gpi/capability#Constraints it states that targets with ZD larger than 45 degrees are not possible in queue.
Data taken in any mode with under CC70 and/or IQ85 requirements are only evaluated based on DIMM/MASS measurements indicating that the data is taken in the required IQ constraint and/or CC constraint.
Observations of Targets fainter or brighter than the magnitude limits are NOT possible in queue.
Update June 16th, 2019
- The GPIES Campaign concluded in semester 2018B.
- Links to relevant Campaign information has moved to the Historical documents section.
- The Duplication policy document has also been updated to reflect the end of the Campaign.
Update February 26th, 2019
- The spot calibration has been updated and the details can be found on the page https://www.gemini.edu/instrumentation/gpi/capability#Spots
Update March 13th, 2020
- GPI is mounted on the telescope. It is still not available for the queue as it is still cooling down. It is expected to be ready for queue by March 17th.
Update February 11th, 2020
- GPI taken off the telescope. Expected to be mounted back March 9th, 2020.
Update December 20th, 2019
- GPI is back on the telescope and ready for observations, if weather permits.
- GPI was taken off for two weeks during the first two weeks of December for the GSAOI Laser run.
Update July 23rd, 2019
- GPI has now been tested out completely after the repairs and is fully functional.
- GPI is planned to be installed back on the telescope on July 29th and thus expected to be operable in queue again on the 31st of July.
Update July 18th, 2019
- The spare high power voltage supply was installed and the tweeter now appears to be responding correctly.
Update June 26th, 2019
- We have identified the failing component and we are working on getting a replacement for the failed board. It is estimated that in a few weeks we will have an update with a firmer timeline on exactly when GPI comes back in queue.
Update June 17th, 2019
- Due to an issue with the tweeter (the high order DM) GPI has been taken off the telescope for further troubleshooting. GPI is thus unavailable for observations until further notice.
Update September 4h, 2018
- GPI was removed from the telescope during the shutdown to repair the IFS computer. As part of the shutdown GPI went through and warmup and cooldown cycle.
- The failed solid state disc in the IFS computer has been replaced and the software reinstalled. Testing showed that the SW reinstall and disc works fine.
Update March 2nd, 2018
- The GPIES Campaign page has been with time accounting and also links to the Gemini Observatory Archive (GOA) up to and including semester 2017B
Update February 28th, 2018
- GPI was removed from the telescope on February 19th in preparation for the Laser commission run, and is expected to be mounted on the telescope again March 5th, 2018.
- GPI available in the 2018B Call for Proposal
Update October 31st, 2017
GPI back on the telescope. With a PV run followed by a GPIES block followed by another PV block, the earliest queue observations would be November 17th.
Update October 13th, 2017
GPI was cooled down and installed on the telescope October 13th
Update October 2nd, 2017
Due to a failure with commercial power and the generator backup the instrument started to warm up on September 21st. The instrument was taken off the telescope and we installed an external pump to reduce the due to the increased to allow a restart of the cooling down without a full warmup cycle. GPI is still in the lab and is warming up.
It is expected to have a full warm-up and pump cycle during the telescope shutdown, October 3rd to 12th, and it is expected that the instrument will be back in operation by October 16th.
Update July 26th, 2017
The AO Computer is now working again as the system and software was restored from backups.
Update July 11th, 2017
GPI will be removed from the telescope Wednesday 12th and the AO Computer will be removed to work on replacing the disk with a recent clone/backup. At the earliest GPI will be back on the telescope no earlier than Wednesday the 19th.
Update July 7th, 2017
GPI has suffered a catastrophic failure of the AO computer disk, and thus unavailable until further notice. We are working hard to replace the disk and the system and we will post information on the situation, by next week.
Update May 29th, 2017
A paper on forward modelling applied to GPI data processing has been added to the section Technical Documents.
Update December 16th, 2016
- The GPIES duplication list has been significantly updated, targets have been removed, added and in some cases polarization mode added to the standard H- observation. The list can be found here.
- GPI back on the telescope and working fine.
Update November 19th, 2016
- NEW FITS keywords added to the headers, both of these keywords are to allow the user to check the validity of the MASS and DIMM values in the header.
MASTSTMP= 6 / MASS time stamp, minutes DIMTSTMP= 0 / DIMM time stamp, minutes
Update October 14th, 2016
- The Sensitivity entry on the Instrument Performance page has been merged with the Contrast information, with updates on Contrast I-magnitude, Contrast Time and Contrast Field Rotation. See the Contrast page for details.
Update October 13th, 2016
- GPI proposals for 2017A consisted of 9% of the total proposed time (excluding Subaru proposals).
- spots have been added at Satellite Spots which is located under the Instrument Performance page.
Update October 11th, 2016
- GPI PPM mask now working, so no restrictions on NRM or other modes.
Update August 9th, 2016
- GPI intervention to repair the PPM mask wheel went well. Testing at warm temperatures worked very well.
- The GPI Calibrations page updated to instruct on how to use the Gemini archive.
- GPI exit wheel fixed in the intervention
Update May 11th, 2016
- The Observing Preparations Page has been updated with details on the now available GS-GPI-library that contains detailed and explained examples of various types of observations.
- On various pages it has been highlighted that observing with GPI in worse conditions than normal is possible, but there is NO guarantee on performance and thus not suitable for high contrast observations like planet detection and discs but rather suitable for binary searches and extended objects like moons and asteroids.
Update February 23rd, 2016
- The Public Data Release page has been updated and reorganized, have look as there is fully reduced available for download.
Update February 19th, 2016
- The pages for Meta modes and the Limiting Magnitudes pages have been updated to reflect the updates relevant to the NRM usage. The NRM mode is still offered in shared risk as we have very limited information on performance and constraints and the data reduction is not fully supported by the GPI pipeline.
Update February 4th, 2016
- A Bad Pixel Map is available on http://www.gemini.edu/sciops/instruments/gpi/data-reduction-and-calibrations/sciops/instruments/gpi/data-reduction-and-calibrations/calibrations#bpm
Update February 4th, 2016
- Weather in the last two months suitable for GPI so significant advance on the GPI queue for 2015B
- Semester 2016A started
- All non-rollover programs from 2015B have been deactivated due to the end of
- PPM mask kept in simulation in APOD_H so no NRM in any band, all other modes possible. Note that in the DIRECT mode a factor of 3 in flux is lost due to APOD_H being in. Simulations and tests show that using the APOD_H in K1 and K2 has a very minimal impact. Note that as the PPM mask is in Simulation the FITS keyword is incorrect for all EXCEPT the H-band.
- GPI comes off the telescope on February 10th, 2016 for a two week period. GPI is expected to be back on the telescope February 22nd and stay on until April 7th. Details on the schedule can be found on http://www.gemini.edu/sciops/metrics/gs2016Aoverview.html
Update November 12th, 2015
- Major update on the GPI OT details page. Detailed description of all the pertinent information to make a GPI observing sequence.
- Active dampers used on sky, updates on performance keep an eye on this webpage for updates
- PPM mask kept in simulation in APOD_H so no NRM in any band nor K2 in any mode observations, all other modes possible. Note that in the DIRECT mode a factor of 3 in flux is lost due to APOD_H being in.
Update September 24th, 2015
- Active dampers are working fine, vibrations down by a factor of 10! With this major milestone achieved it will be possible to advance on the commissioning of the High Order Wavefront Sensor (HOWFS) and using Speckle Updates will be posted with news.
- OMSS entrance shutter replaced
- GPI didn't suffer any damage from the earthquake, aligned and cool in the lab
- PPM mechanism work ongoing
Update September 9th, 2015
While GPI is in the lab repairs will be done on the OMSS entrance shutter, troubleshooting of the PPM mechanism and further work on the active CCR dampers.
Update September 4th, 2015
- The GPIES Campaign target list has been updated with a few changes/additions. Please consult the list at www.gemini.edu/sciops/instruments/gpi/campaign-science/gpi-campaign-target-list before submitting a proposal.
Update August 31st, 2015
- For Semester 2016A the NRM modes and any mode in K2-band are offered without any guarantee for the performance i.e. conditions with no guarantee on the contrast or any other GPI performance indicator.
- GPI is back in operation in the PPM mask is still locked to H-band and thus the GPI Observing Modes are limited to Y/J/H/- only in either spectral or polarization mode. In mode in K1/K2 and any DIRECT mode is not available. We will update this page as soon as we have an update of the status.
Update August 24th, 2015
GPI is now down again operating temperatures after the unplanned warm-up due to the snow storm two weeks ago.
Further testing with the Pupil Plane Mask motor shows that the motor is unreliable as it got stuck again during testing. It is not clear what component is causing the mask to get stuck, and we are continuing the troubleshooting to identify exactly which component in the mask that is causing the mask to get stuck (amplifier, driver card, bearing,. Thus we have in the meantime left the mask in the H-band position. The GPI Observing Modes to Y/J/H/- only in either spectral or polarization mode. In mode in K1/K2 and any DIRECT mode is not available. We will update this page as soon as we have an update of the status.
Update August 6th, 2015
The Pupil Plane Mask suffered a major failure, we have partially recuperate the mask positioning and we have locked the mask in the H-band PPM position. We are continuing diagnosing the failure and in the means that the mask can't be moved and this limits the GPI Observing Modes to H- only in either spectral or polarization mode i.e. H- and H--pol. We will update this page as soon as we have an update of the status, in the assume that for a minimum of two weeks from this date this limit will be in place.
The attempt to install the active vibration dampers in July failed due to an electronics problem and we are waiting for replacement controller boards for the active dampers to attempt the active damper install again. It is expected that the dampers will significantly decrease the 60Hz and multiples vibrations from the CCR's in GPI. This in turn should yield a better contrast on sky, but details will be forthcoming after a successful install and testing on this is uncertain as priority is the repair and diagnosis of the PPM mask.
The new electronics were successfully installed and we now can read out the AOWFS (OIWFS) faster allowing to read an additional row and thus allowing a better control. Bias level drifts were further stabilized by the install of new fans in the cabinet.
The trouble shooting of the 37Hz was successful and we have not seen any 37Hz vibrations in GPI after the fix.
GPI was severely impacted by the poor weather in 2015A semester, as was worse than average and GPI was scheduled in blocks. We will be addressing this by more extensive cross training of observers so that GPI can be fully integrated into the queue operations so that we don't need to schedule GPI in blocks.
- The "Overheads" page has been updated to reflect the now correct detector overheads and half-wave plate changes. It should be noted that acquisition times for non- modes have been decreased to 10minutes from the standard 15minutes. Note that this change will be implemented in a later OT release.
- The pages on "optical throughput", "observing modes", and the "Phase II checks" have been updated so make it clear that the K2 throughput is significantly lower than the K1 throughput. the pages have been updated to make it absolutely clear that we strongly recommend the use of K1 over K2 due to the lower throughput and the higher skies in K2.
- The "Observing Modes" page have been to clarify that the NRM modes have no commissioning data at this time and NRM data are not fully reduced by the pipeline and thus any NRM data are only guaranteed to be taken under the requested conditions.
Update April 20, 2015
There are three major updates:
Reference documents updated with links to data processing publications, very useful for first users of GPI and the GPI pipeline.
- The Phase II checking pages have been used, very useful also for PI's preparing proposals for a quick check of feasibility.
- GPI was the 2nd most popular instrument in LLP proposals (including both GN and GS instruments)
- GPI was the third most popular instrument (9%) in regular proposals, a drop from 2015A as it seems that PI's focused on LLP instead of the regular queue
Update January 30, 2015
- The last night of the semester for GPI tonight. Standard queue operations for GPI start again after March 16th when GPI is back on the telescope.
Update November 20, 2014
- The trouble shooting the bias drift issue and we are now in a situation that GPI will be back in queue. The queue window starts Friday November 21st to Monday November 24th.
- The bias drift was caused by a poor cooling flow inside the Electronics cabinet for the OIWFS. The short term fix was to install an additional fan, and the long term is looking into better temperature control/cooling of the particular electronics boards.
Update November 18, 2014
- It was found at the beginning of the last night of the GPIES Campaign (November 12th) that the bias drifts in the AOWFS were much higher than normal and impeding observation with GPI. Since then we have been investigating the issue and it is most probably related to cooling flow issues inside the electronics cabinet for the AOWFS. As soon as we have news the status will be updated and we will then focus on advancing on the GPI queue programs. We recommend any PI's that have targets that have set and/or close to setting to take contact with the Head of Science Operations (Rene Rutten email@example.com) to get approval for adding new targets that have the same science goal as the original targets and thus allowing the completion of their science goals.
Update November 12, 2014
- The exposure times have been modified and updated on http://www.gemini.edu/sciops/instruments/gpi/instrument-performance/exposure-times. We will update the exposure times in the programs in the queue for 2014B, so no intervention is needed from the PI's.
Update November 5, 2014
- GPI grows in popularity among the Gemini community in the 2015A semester.
- GPI proposals consisted of 16% of the total time requested at Gemini South, up by 5% from last semester.
Update November 4, 2014
On behalf of the GPI data analysis team I am pleased to announce the release of GPI Data Reduction Pipeline version 1.2. Available now from http://docs.planetimager.org/pipeline/. This release provides updates, bug fixes, improved documentation, and new functionality for 2014B and beyond. In particular:
- Improved primitives for flexure correction, ADR compensation, photometric calibration, wavelength calibration, satellite spot photometry and astrometry, and more
- Updated instrument characterization for pixel scale, rotation angle, and photometric zero points
- Recipe generation enhancements and substantial speedups to the Data Parser
- Documentation new tutorials on reducing your own GPI data and calibration
- bunch more.
Detailed release notes and credit to contributors are available at http://docs.planetimager.org/pipeline/installation/relnotes.html
Update November 3, 2014
An email was sent out yesterday to the GPI PI's informing them about the progress of GPI programs in the queue.
You may have seen that we have not been observing your GPI programs, this is not due to technical problems and/or problems with your observations. We have had GPI scheduled in queue blocks this semester, unfortunately the weather has not been cooperative with the expected conditions to observe with GPI. We are continuing with your programs in the queue and we hope that the weather will improve and we can progress with the GPI queue programs.
Update November 2, 2014
Summary of accepted programs in 2014B for GPI.
- A total of 72h of Band 1 and Band programs was accepted in 2014B.
- The time is spread over 12 which 6 are Band 1 programs and 6 are Band 2. Programs are spread over most of the international partners, 2 Australian, 2 Canadian, 4 Chilean and 4 US programs.
- The GPI proposals made up 11% of the total requested time and the over subscription factor for GPI was 3.7
Update May 12, 2014
Detailed instructions on how to obtain the GPI Pipeline and baseline Calibration files from the GSA are available here.
Update May 7, 2014
GPI now listed as a separate instrument at the GSA archive. To query the calibration the complete catalog and select "GPI Standard Query" on the GSA Complete Catalog page.
A bugfix release for the pipeline (v1.1.1) is now available now from http://docs.planetimager.org/pipeline/.
The required acknowledgement for publications using GPI data is given at http://sciopsedit.gemini.edu/sciops/data-and-results/acknowledging-gemini/. The "Public Data" and "Early Science" pages also include this text, but please refer to the previous link as it also includes the standard in LaTeX and other formats.
Update May 5, 2014
- On behalf of the GPI data analysis team I am pleased to announce the release of GPI Data Reduction Pipeline version 1.1. Available now from http://docs.planetimager.org/pipeline/. This release includes numerous enhancements and minor bug fixes from the Feb, March, and April commissioning runs. Detailed release notes and credit to contributors are at http://docs.planetimager.org/pipeline/installation/relnotes.html
Update May 2, 2014
- During the 6 nights dedicated to the GPI Early Science run, 4 nights were granted with and clear sky with a very good seeing. The two last nights were unfortunately plagued with bad weather clouds, winds and extremely poor seeing.
- Out of the 16 accepted programs for the GPI Early Science run, 13 completed, 1 partially completed, 1 attempted and 1 not observed.
- The run was in general very successful and the Gemini GPI team is satisfied by the GPI performance during this observing run.
Update March 24, 2014
- The Early Science programs have been selected and the PI's contacted and they now can work on the Phase II for the observations.
- The second commissioning run started on March 20th and has been a great success, and it is expected that the remaining two nights will be working as well. The run has shown that the ADC is working and it is expected that we can use the ADC as part of the standard operations. Major improvements to the OIWFS control loops have been implemented and work is in progress on the data. The polarization modes have been used frequently to obtain more performance data, high contrast data with the standard and direct imaging data has also been taken. NRM observations has also been taken, which will allow performance evaluations of this mode in time for the 15A CfP.
- It is planned that April 20th to 26th will be the allocated nights the Early Science run.
- Currently the time estimates in the OT wrong as it does not properly calculate the times needed to take an exposure. Please check the Overheads node for more information.
Update February 25, 2014
There has been a lot of activity on the GPI side these last few weeks.
- GPI Early Science has been opened with a deadline of February 28th, 2014
- WWW pages a major overhaul:
- Instrument Performance pages a major update to all
- The Campaign target list has been updated
- The public data more than a dozen targets that can be downloaded. The data sets include both raw and reduced data.
- The remediation was a success, it included:
- Replacement of two motors (pupil and prism/ slide)
- Installation of an IFS baffle that removed the stray light seen on sky flats and GCAL flats
- Software fixes that among other things decreased the overheads in readouts by 4 seconds
- Most importantly a synchronization of the CCR's that decreased the measured vibrations by at least a factor of two
Update September 13, 2013
- GPI is currently in the Acceptance testing stage, which is expected to conclude with the Acceptance review in late May 2013. Assuming the current schedule, the instrument will be delivered to Gemini South in middle 2013. GPI will be used to carry out large Campaign as well as standard science via the TAC system. It is expected that the earliest science observing will be in semester 2014A pending results from Acceptance Testing and Commissioning.
Update September 13, 2013
- The approved GPI campaign program is led by Bruce Macintosh (Lawrence Livermore National Laboratory) and will conduct a comprehensive survey to yield a robust census of extrasolar giant planets, for more information visit this link.
Update August 16, 2011
- GPI is currently under development. Assuming the current schedule, the instrument will be delivered to Gemini South in middle 2012. GPI will be used to carry out large Campaign as well as standard science via the TAC system. It is expected that the earliest science observing will be in semester 2013A pending results from Acceptance Testing and Commissioning.
Update February 27, 2013
- GPI testing at Cerro Pachon have reached a major milestone yesterday by taking images of the calibration light source with the science camera. The final contrast curve is very similar to those obtained in Santa Cruz. Today we ran the Speckle Nulling an improvement of a factor 2 to 3 Progress have been according to the telescope integration plan and it is expected that GPI will go onto the flexure rig in the last week of September.